OpenStudy (poke):

A car's velocity after each consecutive second of motion is 2m/s , 4 m/s ,6m/s . This evidence that the car is A. Moving with a constant speed B. Moving with a constant acceleration C. Moving with a constant velocity

OpenStudy (poke):

Hey

OpenStudy (jiteshmeghwal9):

at t=0 , u=2m/s at t=2, u=4 m/s at t==3, u =6 m/s as u can see the data is following the equation v=u+at. So the body is moving with constant acceleration

OpenStudy (poke):

Which of the following represents an accelerated motion ? Select all that apply ; do not include commas or spaces ... A an object that turns around and changes directions . B. An object that changes its velocity C. An object moving at at constant speed in the same direction. D. An object that remains the at rest E. An object moving in a circle at constant speed . F. None of these

OpenStudy (agent0smith):

I assume they mean non-zero acceleration. If an object is changing speed or velocity, or direction, then it is accelerating.